OpenStudy (anonymous):

Complete the square and write the equation in standard form. Then give the center and radius of the circle. x2 + y2 - 10x - 8y + 29 = 0

OpenStudy (anonymous):

x2-2(5)x+y2-2(4)y+29=0 x2-2(5)x+25-25+y2-2(4)y+16-16+29=0 (x-5)^2+(y-4)^2=12 therefore,center= (5,4) and (radius)^2 = 12 and radius = root(12)=2root(3)

OpenStudy (anonymous):

x2 + y2 - 10x - 8y + 29 = 0 x2 - 10x + y2 - 8y + 29 = 0 x2 - 10x + 25 + y2 - 8y + 16 + 29 = 0 + 25 + 16 (x-5)^2 + (y-4)^2 + 29 = 41 (x-5)^2 + (y-4)^2 = 12 in the form (x-h)^2 + (y-k)^2 = r^2, (h,k) is center and r is radius center is (5,4) and radius is sqrt(12) which is 2sqrt(3)

OpenStudy (anonymous):

medal and fan pls LOL